Avi and Maya are the only two consumers in an economy. Avi’s demand for apples is given by the equation xA(p) = 6 − p, where p is the per unit price of apple. Maya’s demand for apples is given by the equation xB (p) = 10 − p.

(a) Derive the aggregate demand x(p) = xA(p)+xB(p). Draw the aggregate demand where price of apples is on the vertical axis and the aggregate quantity consumed is on the horizontal axis.

(b) Suppose p = 4. Calculate the consumers’ surplus at this price.

(c) Find the elasticity of aggregate demand with respect to price when p = 4.
Answer #1

e) X = - 8 = 10 - X ALP+XB (P) 6-pt lo-p 16-2P 16-20 p= 8-0.59 Price Quantity (6) Cs at 5 a = 16-2x5 16 - 10 cs= 0.5(8-5) (6)